QUESTION 9 A student new to the study of hearing wanted to determine the ability of listeners to determine the difference in the duration of two tones as a means to study temporal discrimination. He asked listeners to discriminate between a tone of 5 mse duration and one of 20 msec duration, and in another condition between a tone of 200 msec duration and one of 400 msec duration. He found that the listeners could detect the difference between the 5 and 20 msec tones better than they could differentiate between the 200 and 400 msec tones. Does this mean that the duration is the only reason that the difference thresholds are different for the short duration tones as compared to the long duration tones? (hint: a short duration stimulus like a click has a spectrum made up of many frequencies). Path: p Words: 0
Answer #1

The frequency is reciprocal of time period, so we can calculate the frequency differences between the waves as, (The frequency will not match with the actual waves frequency, since here the time for observing the wave is different)

yi _ 5-20 0.15 Hz 1 =

\Delta f_2 = \frac{1}{200} - \frac{1}{400} = 2.5 \times 10^{-3} \, Hz

So the person will be able to differentiate between the waves which are farther apart. The duration of the individual tones are also important, since as the duration of the tones decreases, the number of frequencies in the wave increases (Fourier Transform), whereas a long duration tone, has lesser number of frequencies.

One more reason is that the 200 msec and 400 msec duration waves are in frequency region of 1-5 Hz which cannot be heard by Humans.